OpenStudy (anonymous):

if A is a matrix then prove that (A^t)^-1 = (A^-1)^t if A is a matrix then prove that (A^t)^-1 = (A^-1)^t @Mathematics

OpenStudy (jamesj):

Remember first that \( (AB)^T = A^T B^T \) and \( (A^T)^T = A \). Now \( A^T (A^T)^{-1} = I \ \) hence taking the transpose of both sides we have \[ (A^T)^T ((A^T)^{-1})^T = I^T = I \] But \( (A^T)^T = A \) hence \( A ((A^T)^{-1})^T = I \) which means that \[ ((A^T)^{-1})^T = A^{-1} \] Now take the transpose of both sides \[ (A^T)^{-1} = (A^{-1})^T \]

OpenStudy (anonymous):

(AB)^t = B^t A^t isn't it

OpenStudy (jamesj):

Sorry yes. But the rest of the argument holds up. Just flip the order of the matrices where I use that.

OpenStudy (jamesj):

Note to self: do not write mathematics before coffee.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

how can u say A^t (A^t)^-1=I